# A Dynamic Model of Performative Human-ML Collaboration: Theory and Empirical Evidence

### v2

@@ 5 Results
Figure 2: Human Performance Across Treatments. Error bars denote 95% confidence intervals based on standard errors clustered at the user level. Solid bars denote the average performance of the submitted solution, striped bars denote the performance if one picked the higher solution between the submitted solution and the provided ML recommendation.
A total of 1,408 participants completed the study; we removed 119 participants due to forbidden browser reloads or uses of the browser’s back-button, which left 1,289 for the analyses below. See Appendix A.9 for an overview of participants’ demographics. On average, participants’ compensation implied an hourly wage of £12.17 (\$15.22), which is above the UK minimum wage of £11.44. Please see Appendix A.3 for more payment details.
@@
We start by discussing the null results of monetary performance incentives (RQ1). Figure 2(a) shows the results of our monetary incentive experiment. On average, user economic performance without any bonus was 89.7% (light blue bar). None of the bonus alternatives are statistically distinguishable from the control group, nor from each other, and their point estimates are all between 88.6% (for the 20-cent bonus) and 90% (for the 10-cent bonus).
@@
The null effect of monetary incentives is not due to the fact that users did not understand the bonus structure. To test this hypothesis, we can compare the performance of users in the two 10-cent bonus treatments without algorithmic recommendations (third column in Figure 2(a) and first column in Figure 2(b), both yellow). These two treatments only differ by the fact that the one in Figure 2(b) had a comprehension quiz for the bonus structure. The difference in performance between the two treatments is a mere 0.9%, not statistically different from zero ($p=0.268$, based on standard errors clustered at the user level).
@@
RQ2: Because monetary incentives do not change human performance, we test the introduction of ML recommendations with a single bonus structure, the 10-cent bonus. Figure 2(b) presents the results. Focusing on the solid bars, three insights are noteworthy. First, comparing the first two columns (yellow and blue), models with low economic performance seem to lead humans to perform slightly worse than if they were not supported by ML recommendations (89.4% versus 90.9%). This comparison is not statistically significant ($p=0.147$), likely due to low statistical power, but the level difference is not trivial. Despite this, humans do improve performance relative to the algorithmic recommendations (89.4% versus 71.8%, $p=1.8$e-$28$), a result we come back to in Section 5.1. This suggests that people might reduce their effort in solving the problem when they have access to recommendations, but they do not completely eliminate effort (Appendix Figure 18 shows no clear patterns in time spent per problem across treatment conditions). Second, models with better performance lead to increases in human performance, as evidenced by the progressively increasing performance from q1 (72%) to q6 (92%). Third, even if human performance increases with the performance of the ML recommendation, the increments in performance are quantitatively fairly small and sometimes statistically indistinguishable from one another, going from 89.4% when the model’s performance is 72%, to 92.6% when the model’s performance is 92%. [^3] Regression results, controlling for time taken to solve each problem, are presented in Appendix Table 3.
5.1 The Results Within Our Theoretical Framework
Figure 3: Empirical Collaborative Characteristic Function. Confidence intervals are based on standard errors clustered at the participant level.
@@ 5.1 The Results Within Our Theoretical Framework
Figure 3 embeds our empirical results in the theoretical framework presented in Section 3. On the x-axis, we plot the performance of the six ML models deployed in our study. On the y-axis, we plot the performance of the solutions submitted by humans who receive ML recommendations. Each of the points correspond to the six ML treatments of Figure 2(b). We linearly interpolate the estimated points to form an approximation of the collaborative characteristic function $\Delta_{\mathbb{U}}$ (solid blue line). The curve suggests that humans improve on the ML recommendations for ML performance levels between 70% and 92%. The estimated $\delta_{qi}$’s range from 17.5% ($p=1.8$e-$28$) for $q1$, to 0.5% ($p=0.46$) for $q6$. We denote $q6$ a stable point since the human improvement is estimated to be small and statistically indistinguishable from zero. The results imply that, for this portion of the domain, a firm can deploy a model with below-human performance and still converge to a stable point with 92% performance in subsequent deployments.
@@
The $\delta_{qi}$ improvements are always positive (or indistinguishable from zero for q6), but they could have been even larger. Indeed, in this specific setting, as participants add items to the knapsack, in principle, they can easily compare the value of their solution to the value of the ML recommendation (both of which appear at the top of the interface, see Appendix Figure 15). If humans had picked the highest between their solution and the ML recommendation, the collaborative characteristic function would have shifted upward to the dashed green line in Figure 3, and the stable point would have achieved a higher performance than 92%. The discrepancy between the solid and dashed lines increases as the ML model improves, suggesting that even in a straightforward comparison, humans do not follow ML recommendations when it is in their best interest to do so (the difference can also be seen by comparing the solid and striped bars in Figure 2(b)). Appendix Figure 19 decomposes the net effect into two parts. On one hand, as the model performance improves, humans are more likely to follow its recommendations. On the other, when they do not follow the ML recommendation, as the model performance improves, it is much more likely that the submitted solution is inferior compared to the recommendation. Under both solid and dashed collaborative characteristic functions, we can imagine possible collaborative learning paths, $\mathbb{L}_{\Delta_{\mathbb{U}}}$. With this shape of $\Delta_{\mathbb{U}}$, the deployment decision is simple: all collaborative learning paths will eventually reach a stable point at above human performance.
@@ 6 Conclusions
We present a theoretical framework for human-ML collaboration in a dynamic setting where human labels can deviate from the indisputable ground truth. We introduce the collaborative characteristic function, which theoretically links the utility of ML models with respect to the indisputable ground truth, to the utility of humans using those same ML models to support their decisions. The collaborative characteristic function  allows for multiple collaborative learning paths, depending on the utility of the initially deployed ML model. Each of the collaborative learning paths characterizes a possible ML deployment strategy and its ensuing dynamic learning process. We theoretically show conditions under which this dynamic system reaches a stable point through dynamic utility improvement or deterioration. We then present the empirical results of a large user study, which allows us to estimate points on the collaborative characteristic function of the knapsack problem. For ML models of performance between 72% and 92%, our empirical results suggest that the collaborative characteristic function lies above the 45-degree line. Any collaborative learning path starting at utility between 72% and 92% will thus converge to a stable point with utility around 92%. We explore two factors that can shift the collaborative characteristic function. We find that monetary incentives do not seem to affect human performance. However, we find that wherever applicable, a simple post-processing step that picks the best among available solutions (as is possible for the knapsack problem) can substantially shift the collaborative characteristic function upward, leading to stable equilibria of higher utility.
@@
Our work has a number of limitations. On the theoretical side, our collaborative learning paths assume that the firm is able to perfectly replicate human+ML performance in future ML models. Appendix A.7 discusses stability when learning does not exactly replicate previous human+ML performance. On the empirical side, to reduce costs while maintaining statistical power, we only randomized monetary incentives without ML recommendations, and we randomized the quality of ML recommendations while fixing monetary incentives. Studying the interaction of monetary incentives and ML performance is an important extension. The null result of monetary incentives should be interpreted within our context. First, the study participants received payments above minimum wage, and we only tested different levels of linear performance bonuses. It would be valuable to extend our work to evaluate the extent to which alternative base payments or non-linear bonuses may induce different levels of quality and effort by participants and thus collaborative characteristic functions of varying shapes. Our approximation of $\Delta_{\mathbb{U}}$ for the knapsack problem is naturally incomplete since we did not test every possible level of model performance. However, the six points of the curve that we empirically estimate make us fairly comfortable that a linear interpolation is reasonable, at least for model performances between our minimum and maximum.
@@
Future work could investigate the properties of $\Delta_{\mathbb{U}}$ that guarantee a unique optimal stable point, both theoretically and empirically. Provided that researchers have access to the indisputable ground truth, further empirical investigations of collaborative characteristic functions could also shed light on the shape of those functions for practically relevant tasks such as medical diagnoses or hiring decisions. Future work should also discuss fairness aspects of this framework, e.g., whether or not fair stable points exist and how a firm can reach them. More generally, we hope this work generates more interest in studying settings where ML deployments lead to changes in the data generating process, which have broad managerial and practical applications.
@@ A.7 Perfect vs Imperfect learning
In this section we discuss what changes if we loosen the assumption $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We call this assumption the "perfect learner" assumption because the firm perfectly learns the human labels from epoch $t-1$ with a model in epoch $t$. In the following, we consider an imperfect learner such that $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma$. For now, we limit our discussion to the cases of Proposition 1 1 and Proposition 2 2.
In the case of collaborative improvement ($\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$) the human will improve on any model that the firm can deploy. However, if $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma\Rightarrow(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))<0$ then our error $\sigma$ is larger than what we gain by putting a human in the loop. If the above statement is true for all $M_{t}$, then the imperfection has created collaborative harm and we arrived in the case of Proposition 2. However, this means that this would lead to a stable point. If we consider a second scenario where $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma\Rightarrow(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))>0$ for all $M_{t}$. Then we are still in the scenario of collaborative improvement which means that we will reach a stable point. We can think about this when looking at Figure 1. An imperfect learner which is still a collaborative improvement setting is effectively like tilting the green dashed lines with a negative slope. We would still reach point 6-d but would require more iterations than with perfect learning.
In summary, this is an equally empirical question as it is for a perfect learner. The question is much humans improve the system’s performance and in the case of an imperfect learner how much of that improvement gets "eaten up" by learning errors.
@@ A.8 Model Training
Learning to solve the knapsack problem is a research area for itself, however for the small, one-dimensional case of our experiment, it is possible on consumer hardware. We only train models for knapsack instances with 18 items. As input features we concatenate weights $w_{1},...,w_{18}$, values $v_{1},...,v_{18}$, the weight constraint $W$, the sum of the weights and the sum of the values. Thus, our input dimension is 39. Our goal was to train models with a broad spectrum of economic performances, not to solve the knapsack problem perfectly. We added 5 fully connected layers, 4 of them with ReLU activation functions. We use torch.Sigmoid() for our outputs. The output dimension was 18 and the output values in each index can be interpreted as the likelihood that the item belongs to a solution or not. For more details on the architecture, see our code.

### v3

@@ 5 Results
Figure 2: Economic Performance Across Treatments. Error bars denote 95% confidence intervals based on standard errors clustered at the user level. Solid bars denote the average economic performance of the submitted solution, striped bars denote the performance if one picked the higher solution between the submitted solution and the provided ML recommendation. Appendix Figure 4 replicates the analysis using optimality as a measure of utility.
@@
We start by discussing the null results of monetary performance incentives (RQ1). Figure 2(a) shows the results. On average, user economic performance without any bonus is 89.7% (light blue bar). None of the bonus alternatives are statistically distinguishable from the control group, nor from each other, and their point estimates are all between 88.6% (for the 20-cent bonus) and 90% (for the 10-cent bonus).
@@
The null effect of monetary incentives is not due to the fact that users did not understand the bonus structure. To test this hypothesis, we can compare the performance of users in the two 10-cent bonus treatments without algorithmic recommendations (third column in Figure 2(a) and first column in Figure 2(b), both yellow). These two treatments only differ by the fact that the one in Figure 2(b) had a comprehension quiz for the bonus structure. The difference in performance between the two treatments is a mere 0.9%, not statistically different from zero ($p=0.268$, based on standard errors clustered at the user level). If we assume that the effect of monetary incentives without ML support is greater than or equal to their effect with ML support, these results imply that monetary incentives are unlikely to shift the collaborative characteristic function.
@@
RQ2: We test the introduction of ML recommendations with a single bonus structure, the 10-cent bonus. Figure 2(b) presents the results. Focusing on the solid bars, three insights are noteworthy. First, comparing the first two columns (yellow and blue), models with low economic performance seem to lead humans to perform slightly worse than if they were not supported by ML recommendations (89.4% versus 90.9%). This comparison is not statistically significant ($p=0.147$), likely due to low statistical power, but the level difference is not trivial (especially when looking at optimality as a measure of utility in Appendix Figure 4). Despite this, humans’ utility does improve relative to the algorithmic recommendations (89.4% versus 71.8% in the q1 treatment, $p=1.8$e-$28$). Second, models with better economic performance lead to increases in human performance, as evidenced by the progressively increasing economic performance from q1 to q6. Third, even if human performance increases with the performance of the ML recommendation, the increments in performance are quantitatively fairly small and sometimes statistically indistinguishable from one another, going from 89.4% when the model’s performance is 72%, to 92.6% when the model’s performance is 92%. [^3] Regression results, controlling for time taken to solve each problem, are presented in Appendix Table 3. To evaluate whether the results are at least in part due to users changing their effort level, Appendix Figure 21 plots time spent on each problem across treatment conditions, and shows no clear patterns. Additionally, Appendix Figure 10 shows similarly high reported effort levels across ML and non-ML conditions. [^4] Appendix Figure 10 highlights an interesting contrast between users with and without ML recommendations. Indeed, participants without ML stated that they would have exerted less effort if they had been given ML recommendations. In contrast, the majority of those who received ML recommendations believed they would have exerted similar effort even without ML.
5.1 Approximation of the collaborative characteristic function
Figure 3: Empirical approximations of collaborative characteristic functions for two utility functions: economic performance (left) and optimality (right). Error bars represent confidence intervals based on participant-level clustered standard errors. Significance levels for the estimates of $\delta_{qi}$ are based on t-tests against the null $\delta_{qi}=0$. ***: $p<0.001$
@@ 5.1 Approximation of the collaborative characteristic function
Figure 3 embeds our empirical results in the framework presented in Section 3. On the x-axis, we plot the economic performance of the six ML models deployed in our study. On the y-axis, we plot the performance of the solutions submitted by humans who receive ML recommendations: economic performance on the left plot and optimality on the right plot. Each of the points correspond to the six ML treatments of Figure 2(b). We linearly interpolate the estimated points to form an approximation of the collaborative characteristic function $\Delta_{\mathbb{U}}$ (solid blue line). Looking at the left plot, in this approximation of a collaborative characteristic function, humans improve on the ML recommendations for ML performance levels between 70% and 92%. The estimated $\delta_{qi}$’s range from 17.5% ($p=1.8$e-$28$) for $q1$, to 0.5% ($p=0.46$) for $q6$. We denote $q6$ a stable point since the human improvement is estimated to be small and statistically indistinguishable from zero. The results imply that, for this portion of the domain, a firm could deploy a model with below-human performance and still converge to a stable point with 92% performance in subsequent deployments. The insights from the right plot are qualitatively similar, although there is no stable point in the portion of the domain that we explored.
@@
An adjustment to the solution selection method allows us to simulate an additional collaborative characteristic function. Indeed, in this specific setting, as participants add items to the knapsack, in principle, they can easily compare the value of their solution to the value of the ML recommendation (both of which appear at the top of the interface, see Appendix Figure 18). If humans had picked the highest between their solution and the ML recommendation, the collaborative characteristic function would have shifted upward to the dashed green line in Figure 3, and the stable point would have achieved even higher performance. The discrepancy between the solid and dashed lines increases as the ML model improves, suggesting that even in a straightforward comparison, humans do not follow ML recommendations when it is in their best financial interest to do so (the difference can also be seen by comparing the solid and striped bars in Figure 2(b)). Appendix Figure 22 decomposes the net effect into two parts. On one hand, as the model performance improves, humans are more likely to follow its recommendations. On the other, when they do not follow the ML recommendation, as the model performance improves, it is much more likely that the submitted solution is inferior compared to the recommendation. Under both solid and dashed collaborative characteristic functions, we can imagine possible collaborative learning paths, $\mathbb{L}_{\Delta_{\mathbb{U}}}$. With this shape of $\Delta_{\mathbb{U}}$, the deployment decision is simple: all collaborative learning paths will eventually reach a stable point at above human performance.
@@ 6 Conclusions
We present a theoretical framework for human-ML collaboration in a dynamic setting where human labels can deviate from the indisputable ground truth. We introduce the collaborative characteristic function, which theoretically links the utility of ML models with respect to the indisputable ground truth, to the utility of humans using those same ML models to support their decisions. The collaborative characteristic function  allows for multiple collaborative learning paths, depending on the utility of the initially deployed ML model. Each of the collaborative learning paths characterizes a possible ML deployment strategy and its ensuing dynamic learning process. We theoretically show conditions under which this dynamic system reaches a stable point through dynamic utility improvement or deterioration. We then present the empirical results of a large user study, which allows us to approximate collaborative characteristic functions of the knapsack problem. For ML models of economic performance between 72% and 92%, our empirical approximations of collaborative characteristic functions all lie above the 45-degree line. Any collaborative learning path starting at utility between 72% and 92% will thus likely converge to a stable point with utility around 92%. We explore two factors that can shift the collaborative characteristic function. We find that monetary incentives do not seem to affect human performance. However, we find that wherever applicable, a simple post-processing step that picks the best among available solutions (as is possible for the knapsack problem) can substantially shift the collaborative characteristic function upward, leading to stable equilibria of higher utility.
@@
Our work has a number of limitations. On the theoretical side, our collaborative learning paths assume that the firm is able to perfectly replicate human+ML performance in future ML models. Appendix A.7 discusses stability when learning does not exactly replicate previous human+ML performance. However, since this assumption will likely not hold in the real world, imperfect learning may require more iterations than perfect learning, so more empirical studies are required to explore the speed of model convergence. On the empirical side, to reduce costs while maintaining statistical power, we only randomized monetary incentives without ML recommendations, and we randomized the quality of ML recommendations while fixing monetary incentives. Studying the interaction of monetary incentives and ML performance is an important extension. The null result of monetary incentives should be interpreted within our context. Specifically, the study participants received payments above minimum wage, and we only tested different levels of linear performance bonuses. It would be valuable to extend our work to evaluate the extent to which alternative base payments or non-linear bonuses may induce different levels of quality and effort by participants and thus collaborative characteristic functions of varying shapes.
Our approximation of $\Delta_{\mathbb{U}}$ for the knapsack problem is naturally incomplete for two main reasons. First, we use prediction models trained on synthetic data to approximate the collaborative characteristic function. Second, we did not test every possible level of model performance to fully draw the collaborative characteristic function. It is unlikely that these models and their linear interpolation would lead to the same performative trajectories as models trained on human feedback. We see this as a first proof of concept of collaborative characteristic functions, but much more work is needed to estimate these functions in real-world settings.
@@
Future work could investigate the properties of $\Delta_{\mathbb{U}}$ that guarantee a unique optimal stable point, both theoretically and empirically. Provided that researchers have access to the indisputable ground truth, realistic empirical investigations of collaborative characteristic functions are crucial to shed light on the shape of those functions for practically relevant tasks such as medical diagnoses or hiring decisions. Future work should also discuss fairness aspects of this framework, e.g., whether or not fair stable points exist and how a firm can reach them. More generally, we hope this work generates more interest in studying settings where ML deployments lead to changes in the data generating process, which have broad managerial and practical applications.
@@ A.4 Analysis with Optimality
(a) Different bonus incentives.
(b) Different ML recommendations.
Figure 4: Optimality Across Treatments. Error bars denote 95% confidence intervals based on standard errors clustered at the user level. Solid bars denote the average optimality of the submitted solution, striped bars denote the optimality if one picked the best solution between the submitted solution and the provided ML recommendation.
@@ A.7 Perfect vs Imperfect learning
In this section, we discuss what changes if we loosen the assumption $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We call this assumption the "perfect learner" assumption because the firm perfectly learns the human labels from epoch $t-1$ with a model in epoch $t$. In the following, we consider an imperfect learner such that $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma$.
Figure 1 helps illustrate the relaxation of the assumption. An imperfect learner effectively amounts to tilting the 45-degree line upward or downward. The tilt is upward if imperfect learning leads the ML model to have lower performance with respect to the indisputable ground truth compared to the human (i.e., the slope of the straight line is higher than 1). The tilt is downward if imperfect learning leads the ML model to have higher performance with respect to the ground truth (i.e., the slope of the straight line is lower than 1).
It is straightforward to extend Proposition 1 and Proposition 2 to the case of imperfect learning. In the case of collaborative improvement ($\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$), the human will improve on any model that the firm can deploy. However, if imperfect learning leads to $(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))<0$ then the performance gain from the human effort does not fully transfer to the ML model. If the above statement is true for all $M_{t}$, then the imperfection creates collaborative harm, which is the case covered in Proposition 2. However, this would still lead to a stable point. The alternative scenario where $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma\Rightarrow(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))>0$ for all $M_{t}$, is still a scenario of collaborative improvement, which means that we will again reach a stable point.
In summary, imperfect and perfect learners are analogous. In both cases, the crucial question is how much humans improve the system’s performance. For the case of an imperfect learner, an additional empirical question is how much of the human improvement transfers to the ML model.
@@ A.8 Model Training
Figure 6: Distribution of mean model performances trained on human data ex-post to verify that we picked models with reasonable performances. Vertical lines indicate the economic performances of models trained on synthetic data, which were chosen to approximate the collaborative characteristic functionof the task.
Learning to solve the knapsack problem is a research area for itself, however for the small, one-dimensional case of our experiment, it is possible on consumer hardware. We only train models for knapsack instances with 18 items. As input features we concatenate weights $w_{1},...,w_{18}$, values $v_{1},...,v_{18}$, the weight constraint $W$, the sum of the weights and the sum of the values. Thus, our input dimension is 39. Our goal was to train models with a broad spectrum of economic performances, not to solve the knapsack problem perfectly. We added 5 fully connected layers, 4 of them with ReLU activation functions. We use torch.Sigmoid() for our outputs. The output dimension was 18 and the output values in each index can be interpreted as the likelihood that the item belongs to a solution or not. For more details on the architecture, see our code. In summary, all models had dimensions in order of layers: (39,90), (90,550), (550,90), (90,84), (84,18).
@@ A.9 Overview statistics
Figure 10: Task Difficulty, Effort, and Hypothetical Effort with and without ML recommendations. Perceived difficulty of the task does not vary between participants with and without ML recommendations. Participants without ML recommendation expect to invest less effort if they would have had a ML recommendation, while participants with ML recommendations expected to invest the same effort without ML.